The product of 2 consecutive EVEN numbers is 528 find the numbers
Answer by stanbon(57361)

The product of 2 consecutive EVEN numbers is 528 find the numbers
---
1st: 2x
2nd: 2x+2
----
Equation:
2x(2x+2) = 528
4x(x+1) = 528
x(x+1) = 132
x^2+x-132 = 0
(x-11)(x+12) = 0
---
If x = 11; 2x = 22 and 2x+2 = 24
----
If x = -12; 2x = -24 and 2x+2 = -22
===========================================
cheers,
Stan H.
================